Make use of binomial theorem to expand binomial expression & show the range of values of x for which expansion is valid? 1. (2-x)^-2 and 2. (1-x)^1/2

^ = to power of
1
Expert's answer
2020-04-27T19:57:23-0400

The binomial expansion of

"(1+x)^n=1+nx+\frac{n(n-1)}{2!}x^2+...."

For "(1-x)^{\frac12}"

Put "x=-x" and "n=\frac{1}2"

"(1-x)^\frac12=1-\frac{x}{2}+\frac{\frac12(\frac{-1}{2})}{2!}x^2+...."

"-1<-x<1"

"-1<x<1"

For "(2-x)^{-2}=2^{-2}(1-\frac x 2)^{-2}"

"(2-x)^{-2}=2^{-2}[1+x+\frac{3x^2}{4}+....]"

"-1<-(\frac{x}{2})<1"

"-2<x<2"